You are here: Foswiki>AdaptiveOptics Web>Logs>Log_20111130 (01 Dec 2011, GBrusa)Edit Attach

INAF Pisces observing, 2011 November 29-30

Observer: C. Arcidiacono, G. Brusa, K. Boutsia, M. Fumana
Telescope Operator: G. Bechetti
Support Astronomer:
Instrument: Pisces

Summary.

Extreme difficulty in operating the system caused by seeing and software issues. Here a short list:
a) pupil re-rotator does not track the instrument rotator
b) auto-gain routine causes the shell to rip or at least causes skip frames
c) at times large offsets resulted in not being able to close the loop again
d) not very clear problems related to reading buffers from master diagnostics reported by the aos
e) not very clear problems related to timeout reported by the aos
f) some problems related to the freezing of the CCD39 viewer (Not Live)
g) (at the end of the night) the ao loop would build up very large mirror displacements that caused it to rip for no apparent reason

Finally we have had a 'streak' problem with the guider camera. This caused some confusion and 'going around in circles' until we figure we needed to restart the GCS which reinitialized AzCam. Among the things we did (while fighting this streak problem) was to fix the fact that the astigmatism LUT was not being applied (thanks to Marco). The fact that we have been working without the astigmatism correction all yesterday and most of today could indicate that this is indeed an intermittent problem.


Data Summary

See the science log by Konstantina.Boutsia in ObsInaf20111130.

Details

1:14 Preset IC 10 preset failed because camera lens too low but WFS Hardware GUI says 8.90 - 9.40
  Nov 30 01:14:28.875 aos.AcquireReffailed - AcquireRef[9] failed - RETRY: Camera lens temperature too low. T = 5.80, minimum allowed is 6.00 (0) NO_ERROR

1:15 turning off treehouse fans

1:32 preset on IC 10 and now we got maximum skip frames reached, stopping WFS and, found that the script did not wait for the collimation (active optics) to complete.

1:35 preset on IC10 with wait for collimate

1:38 did not converge after several iterations, continue...

1:40 problem with the script, there was an collimate before the preset.. the error in the script caused the shell to apply a large amount of focus which was then corrected by the active optics .... in short do not

1:45 preset on IC 10

1:49 found star 12.7

1:50 the loop was not working because the pupil rotator was not tracking, stop WFS arbitrator process and then start WFS arbitrator on the WFS Control GUI in the telescope tracking re-rotator ON.

1:53 connection to from AOS to Arbitrator lost, AOS thinks WFS is off... so turned OFF from AOS GUI and after the startup now we don't get any errors like AOS.variable name trying to update undefined variable.

1:58 COMMENT: when we send a preset we should check that the shell is set otherwise refuse the command

1:57 preset IC 10

2:01 collimated

2:02 acquire magnitude 12.7

2:05 during the gain optimization the shell received bad commands and the system tried to stopAO unsuccessfully, we had to stop AO from WFS Control GUI and rest and setFlat.

2:06 preset IC 10

2:07 telescope preset completed

2:08 ao failed during auto-gain, the mirror went in safe skip tried StopAO but was not able , eventually the shell went to rest mode and

2:14 preset IC 10

2:19 now the auto-gain was successful it found gTT = 2.70, ho1 = 1.80, ho2 = 0.81.

2:48 when it goes back from offset after 5 minutes of tracking (sky) the star is off by 1-2 arcsec and the acquisition fails, we need to add a quick re-center of the star on the pyramid!

2:51 DIMM gives 0.9" seeing, the guider gives 0.5" but the AO loop is closed, even though the guider star is off-axis it takes some advantage of the AO correction.

3:12 we tried 'fixing' the problem by hand unsuccessfully becuase the software is not able to recover, we need to be able to 'tweak' the telescope pointing.

JMH: After some discussion between John, Guido and Phil, we think that the problem is that nobody is updating the M1 and M2 collimation during the long time in the off-source position. After more thought during the daytime, I think that the return offset should update the collimation and this should not be a problem. The problem could be reduced by doing a faster duty-cycle and spending shorter times in the off-source position.
JMH: See Issue 3835.

5:12 preset on IC 10 but the seeing seems to be very bad and there might be clouds...magnitude 13.6

5:14 preset failed during auto-gain, looking at the guide images you could see a really bad seeing, the FWHM was 2.5 arcsec. DIMM gives 2 arcsec.

5:21 a second preset failed

6:11:04 aos.PresetAOfailed reading shared buffer. Rest shell, Stopped AOS, restarted AOS, stop and start AO Arbitrator

6:22 problems with W, CCD39 not live, AOS sees the hardware status Off.

6:24 solved by turning off and on the hardware.

6:25 preset NGC1560

6:27 preset masterdiagnostic.R.OPTLOOPBUF (-3508) SLOT_NOT_FOUND_ERROR

6:32 stopped and started masterdiagnostics

6:33 preset NGC1560 same problem

6:38 restarted the W, both software and hardware

6:39:59 Command timeout PresetAO[11] failed

6:41 preset NGC1560

6:43 start found 13.6

6:44 closed loop failed (we forgot to stop and start the WFS arbitrator

6:46 preset NGC1560

6:49 closed loop gtt = 2.7 ho1 2.7 ho2 0.96, fwhm on the guider 1.15 arcsec, DIMM 1.2 arcsec

6:51 preset failed

6:53 SetFlatAO

6:54 preset NGC1560

6:55 collimated

6:57 closed loop gtt = 2.7 ho1 2.7 ho2 2.70, fwhm on the guider 1.07 arcsec

7:08 DIMM seeing 0.92 arcsec

7:32 AO loop open because ...

7:36 preset NGC1560 seeing from guider 1.7 arcsec

7:41 StartAO failed Timeout

7:41 preset NGC1560 failed because PrepareAO illegal...

~7:45 preset NGC1560

8:53 DIMM 1 arcsec

9:08 CCD39 Not Live, AO loop open. Restart AO arbitrator does not work. We decided to turn off W hardware and software and restart everything!

9:21 preset NGC1560

9:28 we see a large 'banana' shape

11:21 after two hours (and calls to Marco, Andrew, John) we concluded that:
a) the banana shape was caused mostly by an artifact in the guide camera that was fixed by restarting GCS which reintialized AzCam
b) the astigmatism LUT on the secondary was not being applied and Marco fixed the bug

JMH: See Issue 3834.

11:27 the active preset cannot find the star, Gino is exercising the pointing close to the target.

~11:40 we got going by applying a new flat and clearing active optics

12:00 two attempts at closing the loop failed with the mirror building up a large astigmatism (oriented up down on the screen).

12:08 while sending presets we got an error from the IDL prompt saying "all available logical units are open" (Doug)

12:12 with the support from Alfio we figure that the pupil re-rotator was not tracking (although it was set at an angle different from 0 unlike in the previous cases)

12:13 preset QSO0848 and found star with 14.3

12:19 loop failed

12:23 preset QSO0848

12:26 star with 14.2

12:28 closed loop with gtt = 2.7 gho1 0.61 gho2 0.61

12:29 loop open because of some unknown reason...

13:51 after commanding the power off of the W-unit from the AOS GUI the same gui reported that the unit was on although all the WFS GUIs reported the power off correctly. I had to stop and restart the AOS process to have the ASO GUI reflect the actual status.

-- GuidoBrusa - 30 Nov 2011

-- JohnHill - 30 Nov 2011
Topic revision: r3 - 01 Dec 2011, GBrusa
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback